# A Geometric Algorithm for Tubular Shape Reconstruction from Skeletal Representation

### v2

@@ 3.2 SDF Computation
using the approximated SDF values generates the same result as using true SDF values.
@@ 4.1 Datasets
Ours - 0.1010 1.0000 0.24
ImageCAS Ball Pivoting 0.7006 0.7204 0.8600 0.22
Poisson 0.7548 0.6148 0.9906 0.76
Ours 0.8062 0.5642 0.9996 0.22
ATM22 Ball Pivoting 0.3620 1.2396 0.7040 2.11
Poisson 0.6818 0.7859 0.9957 1.89
Ours 0.7289 0.4325 0.9995 0.93
PARSE2022 Ball Pivoting 0.3282 0.8912 0.7527 8.50
Poisson 0.7178 0.6415 0.9915 4.87
Ours 0.7406 0.4194 0.9988 3.21
DeepVesselNet Ball Pivoting 0.3404 0.6602 0.7943 73.32
Poisson 0.6084 0.3239 0.9987 32.13
Ours 0.6236 0.3017 0.9999 19.54
@@ 4.3 Qualitative and Quantitative Analysis
Tab. 2 shows the quantitative results. Our method demonstrates high superiority in reconstruction accuracy. The recovered mask and shape by our method maintain the highest consistency with original mask and skeleton. Meanwhile, our method is 2.38$\times$ faster than Ball Pivoting Bernardini et al. (1999) and 1.67$\times$ faster than Screened Poisson Kazhdan and Hoppe (2013), respectively. Qualitative results are shown in Fig. 6. Reconstructed shape by Ball Pivoting Bernardini et al. (1999) have more broken connections and flipped faces, therefore it has the lowest reconstruction accuracy. Screened Poisson Kazhdan and Hoppe (2013) tends to reconstruct thicker partial tubes for the areas with sparse skeletal points. The recovered shapes by our method are smoother and achieve the best reconstruction quality. It’s worth noting that the shapes recovered by our method also have smaller volumes than the ones by Screened Poisson Kazhdan and Hoppe (2013).

### v3

@@ 3.2 SDF Computation
using the approximated SDF values generates the same result as using true SDF values. Note that, providing an algorithm to compute real SDF values is not without merit, as real SDF values are more useful in some tasks Park et al. (2019).
@@ 4.1 Datasets
Ours - 0.1013 1.0000 0.54
Ours (fast) - 0.1010 1.0000 0.24
ImageCAS (1507) Ball Pivoting 0.7007 0.7619 0.6999 0.22
Poisson 0.7548 0.4614 0.9939 0.76
Ours 0.7566 0.5983 0.9994 0.39
Ours (fast) 0.7638 0.5511 0.9997 0.22
ATM22 (2688) Ball Pivoting 0.3619 1.2396 0.7040 2.11
Poisson 0.6805 0.7867 0.9929 1.89
Ours 0.6882 0.4884 0.9992 1.56
Ours (fast) 0.7045 0.4048 0.9997 0.93
PARSE2022 (8885) Ball Pivoting 0.2810 0.8912 0.7527 8.50
Poisson 0.6263 0.6415 0.9915 4.87
Ours 0.6347 0.4883 0.9985 4.63
Ours (fast) 0.6516 0.3773 0.9994 3.21
DeepVesselNet (62519) Ball Pivoting 0.2960 0.6602 0.7943 73.32
Poisson 0.7387 0.3239 0.9987 32.13
Ours 0.7473 0.3394 0.9998 25.96
Ours (fast) 0.7498 0.2819 0.9999 19.54
@@ 4.3 Qualitative and Quantitative Analysis
Tab. 2 shows the quantitative results, where we introduce the accuracy and performance of our reconstruction algorithms using real and approximated sdf values, denoted as ours and ours (fast), respectively . Our methods demonstrate high superiority in reconstruction accuracy. While the mask and shape recovered by ours (fast) maintains the highest consistency with original mask and skeleton, the reconstruction accuracy with real sdf values slightly decreases, possibly due to the precision loss incurred during the computation process. Meanwhile, the proposed methods are generally more efficient. Ours (fast) is 2.38$\times$ faster than Ball Pivoting Bernardini et al. (1999) and 1.67$\times$ faster than Screened Poisson Kazhdan and Hoppe (2013), respectively. It’s worth noting that the shapes recovered by our methods also have much smaller volumes than the ones by Screened Poisson Kazhdan and Hoppe (2013) while maintaining a higher reconstruction quality.
@@ 5 Conclusion
Acknowledgements: This work is supported in part by the Natural Science Foundation of China (Grant 62371270) and Shenzhen Key Laboratory of Ubiquitous Data Enabling (No.ZDSYS20220527171406015).